a body moves from rest with a uniform accelaration and travels 270metre in 3second. find the velocity of body at 10second after the start.

Grade:9

1 Answers

PRANAV KUMAR
6 years ago
u=0m/s
s=270m
t=3seconds
by using s=ut+1/2at square
270=0+1/2a*3square
270=9/2a
a=270*2/9
a=60m/s square
now,
for the whole journey of 10 sec,
u=0m/s
t=10sec
a=60m/sec square(as it is uniform)
by using v=u+at
v=0+60*10
v=600
final velocity of body 10 sec after start is 600m/s....